# Mixed Strategy Nash Equilibrium for Crowd Navigation

### v4

@@ Abstract
Robots navigating in crowded areas should negotiate free space with humans rather than fully controlling collision avoidance, as this can lead to freezing behavior. Game theory provides a framework for the robot to reason about potential cooperation from humans for collision avoidance during path planning. In particular, the mixed strategy Nash equilibrium captures the negotiation behavior under uncertainty, making it well suited for crowd navigation. However, computing the mixed strategy Nash equilibrium is often prohibitively expensive for real-time decision-making. In this paper, we propose an iterative Bayesian update scheme over probability distributions of trajectories. The algorithm simultaneously generates a stochastic plan for the robot and probabilistic predictions of other pedestrians’ paths. We prove that the proposed algorithm is equivalent to solving a mixed strategy game for crowd navigation, and the algorithm guarantees the recovery of the global Nash equilibrium of the game. We name our algorithm Bayes’ Rule Nash Equilibrium (BRNE) and develop a real-time model prediction crowd navigation framework. Since BRNE is not solving a general-purpose mixed strategy Nash equilibrium but a tailored formula specifically for crowd navigation, it can compute the solution in real-time on a low-power embedded computer. We evaluate BRNE in both simulated environments and real-world pedestrian datasets. BRNE consistently outperforms non-learning and learning-based methods regarding safety and navigation efficiency. It also reaches human-level crowd navigation performance in the pedestrian dataset benchmark. Lastly, we demonstrate the practicality of our algorithm with real humans on an untethered quadruped robot with fully onboard perception and computation.
I Introduction
@@
The ability to navigate fluently and safely in human-populated spaces is becoming increasingly crucial for deploying robots in real-world environments. Examples include autonomous driving in populated areas [1], service robots at mass events [2], and industrial robots operating alongside human workers [3]. This problem of safe and efficient robot navigation in human crowds with minimal disruption to humans is often referred to as crowd navigation.
Human behavior studies [4, 5] reveal that humans anticipate each other’s collision avoidance behavior when navigating crowds. Empirical results further indicate that navigation algorithms that predict human cooperation for collision avoidance during planning could improve both safety and navigation efficiency [6]. On the other hand, coupling the prediction of cooperative human behavior with motion planning is different from multi-agent planning, such as through crowd motion simulation [7, 8], since humans do not follow pre-defined action rules. Instead, both the humans and the robot should be modeled as optimal planners whose individual objectives depend on others‘ actions [9]. As a result, the simultaneous prediction of cooperative human behavior and robot motion planning becomes the outcome of assessing the coupled optimal actions of all the agents. This design leads to the application of game theory models in crowd navigation, specifically Nash equilibrium [10] as an optimality criterion for the coupled prediction and planning process.
@@
There are two kinds of strategies that can lead to a Nash equilibrium: pure strategy and mixed strategy. We are interested in the latter. Pure strategy refers to choosing deterministic and specific actions, while mixed strategy refers to sampling actions from a probability distribution of possible pure strategies. Crucially, John Nash proved that not all games have a pure strategy Nash equilibrium, but at least one mixed strategy Nash equilibrium exists for all games [11]. Our interest in mixed strategy Nash equilibrium stems from not only its mathematical rigor but also its non-deterministic nature. Prediction based on pure strategy Nash equilibrium would expect humans to follow the paths exactly as predicted, which is too assertive given the uncertain nature of human behavior. Mixed strategy Nash equilibrium, however, maintains such behavioral uncertainty during the robot’s decision-making. A conceptual comparison of traditional navigation framework, cooperative navigation with pure strategy Nash equilibrium, and with mixed strategy Nash equilibrium is shown in Figure 1.
Despite its potential, applying mixed strategy Nash equilibrium to crowd navigation faces several challenges. First, computing mixed strategy Nash equilibrium is often considered impractical due to its computation burden—even for a 3-agent game, the computation can be close to NP-complete [12]. Second, existing Nash equilibrium frameworks often focus on discrete actions and a finite number of strategies (e.g., poker). However, crowd navigation studies physical agents navigating in a continuous space. Third, the practicality of Nash equilibrium depends on how well the game design matches human behavior, yet there is no consensus on a game-theoretic model for crowd navigation.
In this work, we propose a simple iterative Bayesian update scheme for cooperative crowd navigation with mixed strategies. We formally prove that the algorithm guarantees the recovery of the global Nash equilibrium of a mixed strategy game suitable for crowd navigation. We name our algorithm Bayes’ Rule Nash Equilibrium (BRNE). In the BRNE game, each agent aims to minimize the expected risk of collision with other agents, while also minimizing the deviation of the optimal navigation strategy from a nominal navigation strategy. Furthermore, we propose a sampling-based model predictive crowd navigation framework based on BRNE, with the nominal strategies characterized based on Gaussian processes. The proposed navigation framework has a lower time complexity with respect to the number of agents compared to the state-of-the-art game-theoretic planners and runs in real-time on a laptop CPU and a low-power embedded computer. We further integrate the algorithm into an untethered quadruped robot and demonstrate the algorithm’s practicality with real humans using fully onboard perception and computation.
@@
The rest of the paper is organized as follows: Section II reviews the literature on crowd navigation and discusses the key difference between our method and existing methods. We introduce the iterative Bayesian update scheme and the model predictive crowd navigation framework in Section III. Then we show how the algorithm guarantees the convergence to a global Nash equilibrium of a mixed strategy crowd navigation game in Section IV, with extra safety property analysis. Section V contains the details of the evaluation results and the real-world hardware demonstration. Lastly, we conclude the paper and provide further discussion in Section VI.
II Related work
II-A Early work on robot navigation in crowds
@@
Roboticists have been investigating navigation in human environments since the 1990s. Two landmark studies were the RHINO [13] and MINERVA [14] experiments, where robotics systems were deployed in museums to provide tour-guide to thousands of visitors. Additional work for tour-guide robots was also conducted later, such as Robox [15], Mobot [16], Rackham [17], and CiceRobot [18]. A comprehensive review of the history of crowd navigation can be found in [19].
@@
These works use conventional indoor navigation stacks such as the dynamic window approach [20], where humans are modeled as non-reactive obstacles. While these methods are sufficient when the robot interacts with sparse crowds, the robot’s navigation efficiency is limited because of safety concerns. For example, the robot’s speed is intentionally limited to avoid constant emergency stopping [16].
II-B Prediction-then-planning methods
@@
The limitations of conventional navigation methods in human crowds motivate researchers to develop human-aware navigation methods. One prevalent framework in this category is to predict human motion and plan robot actions to avoid hindering human motion, we name such framework as prediction-then-planning. Human motion can be predicted in the form of cost maps, such as from inverse reinforcement learning [21, 22] or unsupervised learning [23]. Human motion can also be directly predicted in forms of trajectories with deep learning methods, such as using long-short term memory models (LSTM) [24], generative adversarial networks (GAN) [25] or graph-based models [26]. We refer readers to [27] for a comprehensive review of human trajectory prediction. Note that, in [28], the authors report that a simple constant velocity model could outperform deep learning-based trajectory prediction methods, arguing that existing neural network architectures are insufficient to capture interpersonal interactions. In addition, another thread of research aims to develop fast reactive planners around dynamic obstacles, such as through field representations [29] or graph search [30].
@@
Planning on top of human motion prediction also introduces the necessity of uncertainty-aware planning: motion planning methods that consider predictive uncertainty in human motion. In [31], a closed-loop belief update of dynamic obstacles’ states is incorporated into receding horizon planning, which reduces the anticipated obstacle uncertainty and generates less conservative navigation behavior. A confidence-ware motion planning framework is proposed in [32] that maintains a Bayesian belief of the human motion prediction confidence with online observations. The planner can be robust against unexpected human behavior by explicitly modeling prediction confidence. Other works have also investigated robust motion planning with uncertain human motion prediction. In particular, [33] introduces a decoupled framework that combines learning-based trajectory prediction [26] with risk-sensitive model predictive control [34]. Lastly, a flipped approach is introduced in [35] in the context of autonomous driving, where a risk-sensitive trajectory prediction approach is proposed for robust planning.
@@
While exhibiting more compliant navigation behavior alongside humans, decoupled prediction and planning is limited by its core assumption—that the robot’s action will not influence humans’ actions. Failing to account for human reaction could lead to robot actions that surprise humans, who in turn react out of the robot’s expectation, resulting in short oscillatory interaction, often referred to as “reciprocal dance” [36]. Failing to account for human reaction in uncertainty quantification could also lead to over-conservative navigation robot behavior—without incorporating human reaction to belief updating during planning, the predictive uncertainty could lead the robot to consider all viable paths are unsafe and the only safe option is to stay still, a phenomenon often referred to as the “freezing robot problem” [6].
II-C Coupled prediction and planning
@@
Given the limitations of decoupled prediction and planning, it becomes necessary to lift the assumption that the robot does not interfere with human intents. Thus, an alternative framework, named coupled prediction and planning, seeks to simultaneously plan robot actions and predict human actions. The seminal work [37] shows that cooperative collision avoidance (CCA)—where humans and robots collectively make decisions for collision avoidance—is crucial for avoiding artifacts such as the “reciprocal dance” and the “freezing robot problem”. The importance of CCA in dense crowds is experimentally verified in a behavioral study [4].
@@
In [6], an individual’s intent is modeled as a Gaussian process and CCA is modeled as a joint decision-making process by coupling Gaussian processes through a collision avoidance-based likelihood function. The statistical optimality of coupled Gaussian processes is further investigated in [38] and [39]. In [40], a distribution space crowd navigation model is proposed, with agent intent modeled as a distribution of trajectories. While similar to the mixed strategy model in this work, the model in [40] is not a game-theory model but is instead a joint decision-making model similar to [6]. Topology-based abstractions are also used for modeling CCA, such as through braids [41, 42] or topological-invariance [43, 44, 45]. CCA is also studied from the perspective of opinion dynamics [46], with a focus on breaking deadlock situations such as “reciprocal dance” [47].
@@
Another group of works focuses on the implicit modeling of cooperative collision avoidance. Implicit CCA models are often obtained from real-world data or simulated human pedestrians—the availability of data will impact the choice of modeling technique. Given the cost of human data collection, implicit modeling with human data cannot afford directly training robot navigation policies with humans; it uses techniques such as imitation learning [48] or inverse reinforcement learning [49, 50]. Aside from collecting human data, an alternative solution is to simulate cooperative navigation agents using decentralized multi-agent collision avoidance methods, such as [7] and [8]. The benefit of implicit CCA modeling from simulated data is that the navigation policies can be trained directly with the reactive agents using reinforcement learning methods, such as in [51], [52], [53], and [54]. However, the sim-to-real transfer of such policies remains an open challenge, mainly due to the distribution shifts between simulated crowd behavior and real-world human behavior. It is also unclear whether the reinforcement learning policies explicitly predict human cooperation during planning, instead of treating humans merely as dynamic obstacles—the latter case falls into the category of the prediction-then-planning framework.
@@
Note that there is a relevant group of works on social convention-aware navigation. The goal is to take into account social conventions during motion planning, such as people’s tendency to walk in groups [55, 56] or people’s preference for not moving when forming certain geometrical structures [57]. Though conceptually similar, we consider these works to be essentially different from coupled prediction and planning, because the social conventions are defined a priori. As a result, these planners do not necessarily consider humans’ reactions to the robot’s actions.
@@
II-D Game-theoretic planning
@@
Even though game-theoretic planning falls into the category of coupled prediction and planning, we separately discuss this body of work here, given its close connection to our work. The key difference between game-theoretic methods and other coupled prediction and planning methods is that game theory assumes each agent makes individual optimal decisions and provides a rigorous optimality criterion for decision-making—the concept of equilibrium. The most commonly used equilibrium notions are Nash equilibrium [10] and Stackelberg equilibrium [58]. Stackelberg equilibrium, in general, has lower computation complexity since it enforces a leader-follower structure. On the other hand, Nash equilibrium is considered the more rigorous and natural notion for crowd navigation, since it assumes equal status for all agents and all agents make decisions simultaneously.
@@
To apply game theory models, crowd navigation is often formulated as a dynamic game: agents, with states governed by certain dynamics, optimize control policies as pure strategies. One of the pioneering works in this direction is [9], where the interaction between a human-driven vehicle and an autonomous vehicle is modeled as a dynamic game. Since then, dynamic game solvers with better numerical efficiency have been developed. In [59], the authors formulate a general-sum dynamic game for crowd navigation, where the local Nash equilibrium is solved by combining an iterative best response scheme and linear-quadratic regulator. In [60], a fast augmented Lagrangian solver is proposed, which converges locally to the generalized Nash equilibrium of a dynamic game and supports real-time model predictive control for autonomous driving. However, pure strategy Nash equilibrium models expect humans to react exactly as predicted, which is too assertive for interacting with humans. Furthermore, existing pure strategy game-theoretic planners can only approximate a locally optimal solution since computing the Nash equilibrium of a generalized game is computationally intractable [12].
@@
Probabilistic variants of pure strategy Nash equilibrium have been proposed to address the lack of flexibility. In [61], an iterative best response scheme is combined with a model predictive path integral (MPPI) control framework to approximate cooperative stochastic control policies between two autonomous vehicles. However, this approach does not formally guarantee recovering a mixed strategy Nash equilibrium. In [62], a stochastic dynamic game formula is proposed to lift the strict rationality assumption of Nash equilibrium and instead assumes bounded rationality of agents. The authors propose a new notion of equilibrium named Entropic Cost Equilibrium (ECE) and show that ECE is equivalent to the mixed strategy Nash equilibrium of a maximum entropy game. The proposed equilibrium formula also enables inverse inference of interaction policy from observations. In [63] and [64], the multimodality in stochastic games is investigated from the perspective of linear-quadratic games and partially observable Markov decision process (POMDP), where the uncertainty originated from the partially observable objectives of non-ego agents. In [65], an explicit mixed strategy game formula is proposed, allowing agents to simultaneously optimize multiple multi-agent pure strategies, but this formula does not recover the mixed strategy Nash equilibrium.
@@
Lastly, game theory models can be combined with other methods for online adaptation during crowd navigation. For example, Nash equilibrium can be used to infer other agents’ internal states, such as altruism. In [66] the pure strategy Nash equilibrium model is used as an inference model for online estimation of human drivers’ social value orientation, leading to the improved prediction accuracy of human driver actions during the interaction. Similar ideas are also explored in [67], [68], and [3]. Another example of bootstrapping adaptive decision-making with game-theoretic models is [69], where the game-theoretic model provides candidates for contingency planning.
@@
There are two major limitations of existing game-theoretic planning methods for crowd navigation. The first limitation is the high computation cost, which prevents existing methods from being applied to real-time navigation in human crowds. For example, as one of the fastest pure strategy dynamic game solvers, ALGAMES [60] still suffers from a time complexity of $O(M^{3})$ with $M$ being the number of players, thus can only perform real-time inference with no more than 2 human pedestrians. The computation of mixed strategy Nash equilibrium is even more burdensome. In [62], the method is also only demonstrated with no more than 2 human agents. In [65], the computation of mixed strategies is offloaded to a neural network trained offline and this approach is only demonstrated for two-agent interaction. The second limitation is the lack of flexibility in uncertainty representation. In [62], the solution is based on the Linear-Quadratic-Gaussian formula. In [65], the mixed strategy model is limited to a small number of pure strategies, instead of a probability distribution.
@@
III Bayesian update for crowd navigation
In this section, we introduce the iterative Bayesian update scheme for crowd navigation. We will focus on the algorithm description and the model predictive navigation framework, leaving the formal properties of the algorithm, such as the guaranteed convergence to a Nash equilibrium, in Section IV. For the same reason, we leave the introduction of the complete game-theoretic formula in Section IV as well.
III-A Notations and definitions

### v6

@@ Abstract
Robots navigating in crowded areas should negotiate free space with humans rather than fully controlling collision avoidance, as this can lead to freezing behavior. Game theory provides a framework for the robot to reason about potential cooperation from humans for collision avoidance during path planning. In particular, the mixed strategy Nash equilibrium captures the negotiation behavior under uncertainty, making it well suited for crowd navigation. However, computing the mixed strategy Nash equilibrium is often prohibitively expensive for real-time decision-making. In this paper, we propose an iterative Bayesian update scheme over probability distributions of trajectories. The algorithm simultaneously generates a stochastic plan for the robot and probabilistic predictions of other pedestrians’ paths. We prove that the proposed algorithm is equivalent to solving a mixed strategy game for crowd navigation, and the algorithm guarantees the recovery of the global Nash equilibrium of the game. We name our algorithm Bayesian Recursive Nash Equilibrium (BRNE) and develop a real-time model prediction crowd navigation framework. Since BRNE is not solving a general-purpose mixed strategy Nash equilibrium but a tailored formula specifically for crowd navigation, it can compute the solution in real-time on a low-power embedded computer. We evaluate BRNE in both simulated environments and real-world pedestrian datasets. BRNE consistently outperforms non-learning and learning-based methods regarding safety and navigation efficiency. It also reaches human-level crowd navigation performance in the pedestrian dataset benchmark. Lastly, we demonstrate the practicality of our algorithm with real humans on an untethered quadruped robot with fully onboard perception and computation.
keywords:
Crowd navigation; Game theory; Gaussian processes
1 Introduction
@@
The ability to navigate fluently and safely in human-populated spaces is becoming increasingly crucial for deploying robots in real-world environments. Examples include autonomous driving in populated areas (Bai et al., 2015), service robots at mass events (Singh et al., 2021), and industrial robots operating alongside human workers (Bansal et al., 2022). This problem of safe and efficient robot navigation in human crowds with minimal disruption to humans is often referred to as crowd navigation.
Human behavior studies (Murakami et al., 2021; Bacik et al., 2023) reveal that humans anticipate each other’s collision avoidance behavior when navigating crowds. Empirical results further indicate that navigation algorithms that predict human cooperation for collision avoidance during planning could improve both safety and navigation efficiency (Trautman et al., 2015). On the other hand, coupling the prediction of cooperative human behavior with motion planning is different from multi-agent planning, such as through crowd motion simulation (Helbing and Molnár, 1995; van den Berg et al., 2011), since humans do not follow pre-defined action rules. Instead, both the humans and the robot should be modeled as optimal planners whose individual objectives depend on others‘ actions (Sadigh et al., 2018). As a result, the simultaneous prediction of cooperative human behavior and robot motion planning becomes the outcome of assessing the coupled optimal actions of all the agents. This design leads to the application of game theory models in crowd navigation, specifically Nash equilibrium (Nash, 1950) as an optimality criterion for the coupled prediction and planning process.
@@
There are two kinds of strategies that can lead to a Nash equilibrium: pure strategy and mixed strategy. We are interested in the latter. Pure strategy refers to choosing deterministic and specific actions, while mixed strategy refers to sampling actions from a probability distribution of possible pure strategies. Crucially, John Nash proved that not all games have a pure strategy Nash equilibrium, but at least one mixed strategy Nash equilibrium exists for all games (Nash, 1951). Our interest in mixed strategy Nash equilibrium stems from not only its mathematical rigor but also its non-deterministic nature. Prediction based on pure strategy Nash equilibrium would expect humans to follow the paths exactly as predicted, which is too assertive given the uncertain nature of human behavior. Mixed strategy Nash equilibrium, however, maintains such behavioral uncertainty during the robot’s decision-making. A conceptual comparison of traditional navigation framework, cooperative navigation with pure strategy Nash equilibrium, and with mixed strategy Nash equilibrium is shown in Figure 1.
Despite its potential, applying mixed strategy Nash equilibrium to crowd navigation faces several challenges. First, computing mixed strategy Nash equilibrium is often considered impractical due to its computation burden—even for a 3-agent game, the computation can be close to NP-complete (Daskalakis et al., 2009). Second, existing Nash equilibrium frameworks often focus on discrete actions and a finite number of strategies (e.g., poker). However, crowd navigation studies physical agents navigating in a continuous space. Third, the practicality of Nash equilibrium depends on how well the game design matches human behavior, yet there is no consensus on a game-theoretic model for crowd navigation.
In this work, we propose a simple iterative Bayesian update scheme for cooperative crowd navigation with mixed strategies. We formally prove that the algorithm guarantees the recovery of the global Nash equilibrium of a mixed strategy game suitable for crowd navigation. We name our algorithm Bayesian Recursive Nash Equilibrium (BRNE). In the BRNE game, each agent aims to minimize the expected risk of collision with other agents, while also minimizing the deviation of the optimal navigation strategy from a nominal navigation strategy. Furthermore, we propose a sampling-based model predictive crowd navigation framework based on BRNE, with the nominal strategies characterized based on Gaussian processes. The proposed navigation framework has a lower time complexity with respect to the number of agents compared to the state-of-the-art game-theoretic planners and runs in real-time on a laptop CPU and a low-power embedded computer. We further integrate the algorithm into an untethered quadruped robot and demonstrate the algorithm’s practicality with real humans using fully onboard perception and computation.
@@
The rest of the paper is organized as follows: Section 2 reviews the literature on crowd navigation and discusses the key difference between our method and existing methods. We introduce the iterative Bayesian update scheme and the model predictive crowd navigation framework in Section 3. Then we show how the algorithm guarantees the convergence to a global Nash equilibrium of a mixed strategy crowd navigation game in Section 4, with extra safety property analysis. Section 5 contains the details of the evaluation results and the real-world hardware demonstration. Lastly, we conclude the paper and provide further discussion in Section 6.
2 Related work
2.1 Early work on robot navigation in crowds
@@
Roboticists have been investigating navigation in human environments since the 1990s. Two landmark studies were the RHINO (Burgard et al., 1998) and MINERVA (Thrun et al., 2000) experiments, where robotics systems were deployed in museums to provide tour-guide to thousands of visitors. Additional work for tour-guide robots was also conducted later, such as Robox (Siegwart et al., 2003), Mobot (Nourbakhsh et al., 2003), Rackham (Clodic et al., 2006), and CiceRobot (Chella and Macaluso, 2009). A comprehensive review of the history of crowd navigation can be found in (Mavrogiannis et al., 2023b).
@@
These works use conventional indoor navigation stacks such as the dynamic window approach (Fox et al., 1997), where humans are modeled as non-reactive obstacles. While these methods are sufficient when the robot interacts with sparse crowds, the robot’s navigation efficiency is limited because of safety concerns. For example, the robot’s speed is intentionally limited to avoid constant emergency stopping (Nourbakhsh et al., 2003).
2.2 Prediction-then-planning methods
@@
The limitations of conventional navigation methods in human crowds motivate researchers to develop human-aware navigation methods. One prevalent framework in this category is to predict human motion and plan robot actions to avoid hindering human motion, we name such framework as prediction-then-planning. Human motion can be predicted in the form of cost maps, such as from inverse reinforcement learning (Ziebart et al., 2009; Luber et al., 2012) or unsupervised learning (Henry et al., 2010). Human motion can also be directly predicted in forms of trajectories with deep learning methods, such as using long-short term memory models (LSTM) (Alahi et al., 2016), generative adversarial networks (GAN) (Gupta et al., 2018) or graph-based models (Salzmann et al., 2020). We refer readers to Rudenko et al. (2020) for a comprehensive review of human trajectory prediction. Note that, in Schöller et al. (2020), the authors report that a simple constant velocity model could outperform deep learning-based trajectory prediction methods, arguing that existing neural network architectures are insufficient to capture interpersonal interactions. In addition, another thread of research aims to develop fast reactive planners around dynamic obstacles, such as through field representations (Huber et al., 2022) or graph search (Cao et al., 2019).
@@
Planning on top of human motion prediction also introduces the necessity of uncertainty-aware planning: motion planning methods that consider predictive uncertainty in human motion. In Du Toit and Burdick (2012), a closed-loop belief update of dynamic obstacles’ states is incorporated into receding horizon planning, which reduces the anticipated obstacle uncertainty and generates less conservative navigation behavior. A confidence-ware motion planning framework is proposed in Fridovich-Keil et al. (2020a) that maintains a Bayesian belief of the human motion prediction confidence with online observations. The planner can be robust against unexpected human behavior by explicitly modeling prediction confidence. Other works have also investigated robust motion planning with uncertain human motion prediction. In particular, Nishimura et al. (2020) introduces a decoupled framework that combines learning-based trajectory prediction (Salzmann et al., 2020) with risk-sensitive model predictive control (Nishimura and Schwager, 2021). Lastly, a flipped approach is introduced in Nishimura et al. (2023) in the context of autonomous driving, where a risk-sensitive trajectory prediction approach is proposed for robust planning.
@@
While exhibiting more compliant navigation behavior alongside humans, decoupled prediction and planning is limited by its core assumption—that the robot’s action will not influence humans’ actions. Failing to account for human reaction could lead to robot actions that surprise humans, who in turn react out of the robot’s expectation, resulting in short oscillatory interaction, often referred to as “reciprocal dance” (Feurtey, 2000). Failing to account for human reaction in uncertainty quantification could also lead to over-conservative navigation robot behavior—without incorporating human reaction to belief updating during planning, the predictive uncertainty could lead the robot to consider all viable paths are unsafe and the only safe option is to stay still, a phenomenon often referred to as the “freezing robot problem” (Trautman et al., 2015).
2.3 Coupled prediction and planning
@@
Given the limitations of decoupled prediction and planning, it becomes necessary to lift the assumption that the robot does not interfere with human intents. Thus, an alternative framework, named coupled prediction and planning, seeks to simultaneously plan robot actions and predict human actions. The seminal work (Trautman and Krause, 2010) shows that cooperative collision avoidance (CCA)—where humans and robots collectively make decisions for collision avoidance—is crucial for avoiding artifacts such as the “reciprocal dance” and the “freezing robot problem”. The importance of CCA in dense crowds is experimentally verified in a behavioral study (Murakami et al., 2021).
@@
In Trautman et al. (2015), an individual’s intent is modeled as a Gaussian process and CCA is modeled as a joint decision-making process by coupling Gaussian processes through a collision avoidance-based likelihood function. The statistical optimality of coupled Gaussian processes is further investigated in Trautman (2017) and Trautman and Patel (2020). In Sun et al. (2021), a distribution space crowd navigation model is proposed, with agent intent modeled as a distribution of trajectories. While similar to the mixed strategy model in this work, the model in Sun et al. (2021) is not a game-theory model but is instead a joint decision-making model similar to Trautman et al. (2015). Topology-based abstractions are also used for modeling CCA, such as through braids (Mavrogiannis et al., 2017; Mavrogiannis and Knepper, 2019) or topological-invariance (Mavrogiannis et al., 2018; Mavrogiannis and Knepper, 2021; Mavrogiannis et al., 2023a). CCA is also studied from the perspective of opinion dynamics (Bizyaeva et al., 2023), with a focus on breaking deadlock situations such as “reciprocal dance” (Cathcart et al., 2023).
@@
Another group of works focuses on the implicit modeling of cooperative collision avoidance. Implicit CCA models are often obtained from real-world data or simulated human pedestrians—the availability of data will impact the choice of modeling technique. Given the cost of human data collection, implicit modeling with human data cannot afford directly training robot navigation policies with humans; it uses techniques such as imitation learning (Kim and Pineau, 2013) or inverse reinforcement learning (Kim and Pineau, 2016; Kretzschmar et al., 2016). Aside from collecting human data, an alternative solution is to simulate cooperative navigation agents using decentralized multi-agent collision avoidance methods, such as Helbing and Molnár (1995) and van den Berg et al. (2011). The benefit of implicit CCA modeling from simulated data is that the navigation policies can be trained directly with the reactive agents using reinforcement learning methods, such as in Chen et al. (2017), Chen et al. (2019), Liang et al. (2020), and Liu et al. (2021). However, the sim-to-real transfer of such policies remains an open challenge, mainly due to the distribution shifts between simulated crowd behavior and real-world human behavior. It is also unclear whether the reinforcement learning policies explicitly predict human cooperation during planning, instead of treating humans merely as dynamic obstacles—the latter case falls into the category of the prediction-then-planning framework.
@@
Note that there is a relevant group of works on social convention-aware navigation. The goal is to take into account social conventions during motion planning, such as people’s tendency to walk in groups (Wang et al., 2022; Sathyamoorthy et al., 2022) or people’s preference for not moving when forming certain geometrical structures (Sathyamoorthy et al., 2020). Though conceptually similar, we consider these works to be essentially different from coupled prediction and planning, because the social conventions are defined a priori. As a result, these planners do not necessarily consider humans’ reactions to the robot’s actions.
@@
2.4 Game-theoretic planning
@@
Even though game-theoretic planning falls into the category of coupled prediction and planning, we separately discuss this body of work here, given its close connection to our work. The key difference between game-theoretic methods and other coupled prediction and planning methods is that game theory assumes each agent makes individual optimal decisions and provides a rigorous optimality criterion for decision-making—the concept of equilibrium. The most commonly used equilibrium notions are Nash equilibrium (Nash, 1950) and Stackelberg equilibrium (Von Stackelberg, 2011). Stackelberg equilibrium, in general, has lower computation complexity since it enforces a leader-follower structure. On the other hand, Nash equilibrium is considered the more rigorous and natural notion for crowd navigation, since it assumes equal status for all agents and all agents make decisions simultaneously.
@@
To apply game theory models, crowd navigation is often formulated as a dynamic game: agents, with states governed by certain dynamics, optimize control policies as pure strategies. One of the pioneering works in this direction is Sadigh et al. (2018), where the interaction between a human-driven vehicle and an autonomous vehicle is modeled as a dynamic game. Since then, dynamic game solvers with better numerical efficiency have been developed. In Fridovich-Keil et al. (2020b), the authors formulate a general-sum dynamic game for crowd navigation, where the local Nash equilibrium is solved by combining an iterative best response scheme and linear-quadratic regulator. In Le Cleac’h et al. (2022), a fast augmented Lagrangian solver is proposed, which converges locally to the generalized Nash equilibrium of a dynamic game and supports real-time model predictive control for autonomous driving. However, pure strategy Nash equilibrium models expect humans to react exactly as predicted, which is too assertive for interacting with humans. Furthermore, existing pure strategy game-theoretic planners can only approximate a locally optimal solution since computing the Nash equilibrium of a generalized game is computationally intractable Daskalakis et al. (2009).
@@
Probabilistic variants of pure strategy Nash equilibrium have been proposed to address the lack of flexibility. In Williams et al. (2018), an iterative best response scheme is combined with a model predictive path integral (MPPI) control framework to approximate cooperative stochastic control policies between two autonomous vehicles. However, this approach does not formally guarantee recovering a mixed strategy Nash equilibrium. In Mehr et al. (2023), a stochastic dynamic game formula is proposed to lift the strict rationality assumption of Nash equilibrium and instead assumes bounded rationality of agents. The authors propose a new notion of equilibrium named Entropic Cost Equilibrium (ECE) and show that ECE is equivalent to the mixed strategy Nash equilibrium of a maximum entropy game. The proposed equilibrium formula also enables inverse inference of interaction policy from observations. In So et al. (2023) and So et al. (2022), the multimodality in stochastic games is investigated from the perspective of linear-quadratic games and partially observable Markov decision process (POMDP), where the uncertainty originated from the partially observable objectives of non-ego agents. In Peters et al. (2022), an explicit mixed strategy game formula is proposed, allowing agents to simultaneously optimize multiple multi-agent pure strategies, but this formula does not recover the mixed strategy Nash equilibrium.
@@
Lastly, game theory models can be combined with other methods for online adaptation during crowd navigation. For example, Nash equilibrium can be used to infer other agents’ internal states, such as altruism. In Schwarting et al. (2019) the pure strategy Nash equilibrium model is used as an inference model for online estimation of human drivers’ social value orientation, leading to the improved prediction accuracy of human driver actions during the interaction. Similar ideas are also explored in Peters et al. (2020), Le Cleac’h et al. (2021), and Bansal et al. (2022). Another example of bootstrapping adaptive decision-making with game-theoretic models is Peters et al. (2024), where the game-theoretic model provides candidates for contingency planning.
@@
There are two major limitations of existing game-theoretic planning methods for crowd navigation. The first limitation is the high computation cost, which prevents existing methods from being applied to real-time navigation in human crowds. For example, as one of the fastest pure strategy dynamic game solvers, ALGAMES (Le Cleac’h et al., 2022) still suffers from a time complexity of $O(M^{3})$ with $M$ being the number of players, thus can only perform real-time inference with no more than 2 human pedestrians. The computation of mixed strategy Nash equilibrium is even more burdensome. In Mehr et al. (2023), the method is also only demonstrated with no more than 2 human agents. In Peters et al. (2022), the computation of mixed strategies is offloaded to a neural network trained offline and this approach is only demonstrated for two-agent interaction. The second limitation is the lack of flexibility in uncertainty representation. In Mehr et al. (2023), the solution is based on the Linear-Quadratic-Gaussian formula. In Peters et al. (2022), the mixed strategy model is limited to a small number of pure strategies, instead of a probability distribution.
@@
3 Bayesian update for crowd navigation
In this section, we introduce the iterative Bayesian update scheme for crowd navigation. We will focus on the algorithm description and the model predictive navigation framework, leaving the formal properties of the algorithm, such as the guaranteed convergence to a Nash equilibrium, in Section 4. For the same reason, we leave the introduction of the complete game-theoretic formula in Section 4 as well.
3.1 Notations and definitions